### v1

Abstract
Reconstructing full fields from extremely sparse and random measurements is a longstanding ill-posed inverse problem. A powerful framework for addressing such challenges is hierarchical probabilistic modeling, where uncertainty is represented by intermediate variables and resolved through marginalization during inference. Inspired by this principle, we propose Cascaded Sensing (Cas-Sensing), a hierarchical reconstruction framework that integrates an autoencoder–diffusion cascade. First, a neural operator-based functional autoencoder reconstructs the dominant structures of the original field—including large-scale components and geometric boundaries—from arbitrary sparse inputs, serving as an intermediate variable. Then, a conditional diffusion model, trained with a mask-cascade strategy, generates fine-scale details conditioned on these large-scale structures. To further enhance fidelity, measurement consistency is enforced via the manifold constrained gradient based on Bayesian posterior sampling during the generation process. This cascaded pipeline substantially alleviates ill-posedness, delivering accurate and robust reconstructions. Experiments on both simulation and real-world datasets demonstrate that Cas-Sensing generalizes well across varying sensor configurations and geometric boundaries, making it a promising tool for practical deployment in scientific and engineering applications.
keywords:
Field reconstruction; Probabilistic modeling; Cascade pipeline; Functional autoencoder; Diffusion model.
1 Introduction
Despite advances in sensor technology enabling large-scale deployments in physical systems, the number of monitoring sites is often constrained by high setup and maintenance costs. Therefore, spatial field reconstruction from limited local sensor information is a major challenge of complex physical systems, such as astrophysics [akiyama2019first], geophysics [kondrashov2006spatio, carrassi2018data, manohar2018data], atmospheric science [alonso2010novel, mishra2014compressed], and fluid dynamics [vinuesa2023transformative, buzzicotti2023data]. Due to the limited sensors, the reconstruction of multi-scale full-field data from extremely sparse and random observations is inherently an ill-posed problem, where the results often suffers from issues related to existence, uniqueness, and stability [hadamard2014lectures, breiding2023algebraic].
Compressed sensing has been widely applied to address this challenge. It relies on the assumption that signals are sparse in a specific transform domain, and reconstructs full-field data from undersampled observations by solving optimization problems [candes2006robust]. To enhance reconstruction performance, compressed sensing variants combined with deep learning have been proposed, aiming to automatically learn effective sparse representations and reconstruction mappings in a data-driven manner [sun2017intelligent, shi2019image, ni2020deep, xu2022sparse]. While these methods have achieved certain theoretical and experimental progress, several limitations remain. Classic compressed sensing heavily depends on the sparsity prior, leading to performance degradation when dealing with non-sparse structures. Deep learning-based variants, although more flexible, are still often restricted to single-scale, fixed-domain reconstruction and struggle to capture multi-scale information.
Besides the compressed sensing-based methods, many existing methods attempt to directly establish an end-to-end mapping between sparse measurements and the corresponding full-field data [fukami2021global, fan2025vitae, li2025novel, ghazijahani2025spatial]. However, such mappings are typically trained with a specific reorganization form of sparse data (for example, patch or voronoi tessellation) and geometric boundary. When encountering unseen sparse measurements—such as data acquired from a different sensor configuration—or when facing new geometric boundaries, these models often fail to produce accurate reconstructions and require retraining to adapt to the new scenarios. As a result, the applicability of these methods is generally limited to controlled or narrowly defined settings. They often struggle with generalization when applied to complex, real-world environments. Moreover, when the available observations become extremely sparse, the inherent ill-posed nature of the reconstruction problem severely impedes the ability to establish a stable and reliable mapping from sparse inputs to the full-field data [bell1978solutions, candes2011tight, arridge2019solving].
Generative models, such as generative adversarial networks [goodfellow2020generative] and diffusion models [ho2020denoising, song2020score], have shown remarkable capabilities in producing high-quality, diverse data, with broad applications in vision [wang2023dr2, miao2024waveface, zhao2023towards], biomedicine [cao2024high, yu2020medical, nguyen2024sequence], and scientific computing [lienen2306zero, ruhling2023dyffusion, li2024synthetic]. By learning data distributions rather than explicit input–output mappings, these models can generate unseen but statistically consistent samples. Early studies applying generative models to field reconstruction primarily follow two approaches. One line of work learns supervised conditional distributions between full-field low-resolution and high-resolution data pairs [deng2019super, shu2023physics]. However, due to the immense variability of sparse measurement patterns, directly modeling the distribution of full-field data conditioned on sparse observations is practically infeasible, as it is impossible to cover all sparse-to-full field mappings during training. Another approach adopts a reverse diffusion step followed by a projection-based measurement consistency step [choi2021ilvr, chung2022come, kadkhodaie2021stochastic] or a Bayesian sampling step [du2024conditional, li2024learning] after pretraining an unconditional diffusion model to ensure the generated samples match sparse observations. However, this enforced measurement consistency often throws the sample path off the data manifold, leading to incorrect reconstructions [chung2022improving]. Although the manifold constrained gradient [chung2022improving] is proposed to mitigate this problem, under extremely sparse observations, the model still tends to produce random samples due to the overly loose constraints, resulting in non-unique and unreliable reconstructions. Similarly, FunDiff [wang2025fundiff] also relies on incorporating measurement consistency during the sampling stage to enforce conditional generation, and therefore suffers from degraded reconstruction quality under extremely sparse observations. Moreover, it extends diffusion models to function spaces by coupling a function autoencoder with latent diffusion, enabling physics-informed generative modeling across different discretizations. While powerful for capturing global structures, this latent-space generation strategy introduces an information bottleneck: fine-scale details may be lost during encoding and cannot be fully recovered during decoding, which can limit its effectiveness in detail reconstruction from sparse observations.
Cascading pipeline of generative model is an effective way to improve the generation quality in case without strong conditioning information [nichol2021improved, saharia2022image, ho2022cascaded]. In computer vision, this approach involves training a sequence of models at progressively higher resolutions: a base model generates low-resolution samples, which are then refined by super-resolution models [ho2022cascaded]. Compared to training a single high-resolution model, cascading allocates most model capacity to low resolutions, which are critical for overall sample quality, while significantly improving training and inference efficiency. Moreover, individual models can be optimized independently, with architecture choices tailored to each resolution for optimal pipeline performance.
Inspired by the cascading pipeline, the Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade is proposed to achieve accurate and robust multi-scale fields reconstruction from extremely sparse measurements. In Cas-Sensing, by viewing the full-field data as functional data, an neural operator-based functional autoencoder is first trained [bunker2024autoencoders] to extract the latent features of data, thus having the ability of robustly recovering the principal structures (large-scale components with geometric boundaries) from arbitrary sparse measurements. Then a conditional diffusion model is trained with the conditions—the reconstructed principal structures—to only focus on generating the small-scale details which are the difference between ground truth and conditions. To enhance the robustness of diffusion model, we introduce a Mask-Cascade Training (MCT) approach, which applies different random masks during training to generate diverse conditions from a fixed ground-truth field. After training the conditional diffusion model, we employ the Manifold Constrained Gradient (MCG) [chung2022improving] during sampling to ensure that the generated full-field data not only matches the sparse observations at measurement locations, but also remains consistent with the underlying data manifold. The proposed cascade pipeline effectively alleviates the ill-posedness of the original reconstruction problem by progressively recovering multi-scale components in a staged manner, thereby enabling Cas-Sensing to achieve high accuracy, strong generalizability, and enhanced robustness in reconstruction tasks.
The remainder of the paper is organized as follows. In Section 2, we present the problem formulation and describe the proposed methodology in detail, including the functional autoencoder, the conditional diffusion model, and the proposed mask-cascade training approach. Section 3 demonstrates the effectiveness of Cas-Sensing through a series of experiments with simulated data of circular cylinder flow, real-world data of ocean wave height, and real-world data of global ocean temperature. In Section 4, we discuss the advantages and limitations of Cas-Sensing. Finally, Section 5 concludes the paper and outlines directions for future work.
2 Methods
2.1 Problem formulation
The problem of recovering the multi-scale full field $\boldsymbol{u}\in\mathbb{R}^{m}$ from a vector of sparse sensor measurements $\boldsymbol{y}\in\mathbb{R}^{n}$ can be formulated as modeling the conditional distribution $p(\boldsymbol{u}|\boldsymbol{y})$. However, when the observations are excessively sparse, directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$ is ill-posed, as the solution becomes highly non-unique. To address this challenge, we propose Cas-Sensing, a cascaded framework that mitigates the ill-posedness by leveraging the following marginalization rule:
$p(\boldsymbol{u}|\boldsymbol{y})=\int p(\boldsymbol{u},\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}=\int p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}=\mathbb{E}_{\boldsymbol{m}\sim p(\boldsymbol{m}|\boldsymbol{y})}[\,p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,],$ (1)
where $\boldsymbol{m}$ represents the principal structure (the large-scale component) of $\boldsymbol{u}$. This formulation naturally decomposes the task into two subtasks: modeling $p(\boldsymbol{m}|\boldsymbol{y})$ to infer the principal structure from sparse observations, and modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ to reconstrut the full field $\boldsymbol{u}$ by adding high-frequency details to $\boldsymbol{m}$ based on both $\boldsymbol{m}$ and $\boldsymbol{y}$. Since the principal structures of the signal are often sparse, reconstructing $\boldsymbol{m}$ from sparse observations—i.e., modeling $p(\boldsymbol{m}|\boldsymbol{y})$—is generally well-posed. Furthermore, compared to directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$, modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ incorporates an additional condition, constraining the reconstruction of $\boldsymbol{u}$ to remain consistent with the principal structures $\boldsymbol{m}$. Consequently, the degree of ill-posedness is substantially alleviated through this cascaded pipeline.
In the proposed Cas-Sensing framework based on an autoencoder-diffusion cascade, a functional autoencoder is first trained to approximately model $p(\boldsymbol{m}|\boldsymbol{y})$. Then, a conditional diffusion model trained with the proposed mask-cascade training strategy to approximate $\mathbb{E}_{\boldsymbol{m}\sim p(\boldsymbol{m}|\boldsymbol{y})}[\,p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,]$. Specifically, once the autoencoder is trained and fixed, random sparse measurements $\boldsymbol{y}$ are generated from $\boldsymbol{u}$ and mapped through the autoencoder to produce corresponding $\boldsymbol{m}$. The conditional diffusion model is then trained to generate $\boldsymbol{u}$ conditioned on both $\boldsymbol{m}$ and $\boldsymbol{y}$, effectively learning to reconstruct the full field from varying sparse observations.
2.2 Functional autoencoder for principal structure reconstruction
In practical engineering, sensors are often randomly placed, and many physical fields involve complex geometric boundaries, making field reconstruction from unstructured data challenging [chai2020deep]. To address this, we adopt a neural operator-based functional autoencoder [bunker2024autoencoders], which enables coordinate-based input for neural networks[regazzoni2024learning, krishnapriyan2023learning]. The functional autoencoder extracts latent features from unstructured data and robustly reconstruct the principal structure from random and sparse inputs.
We model the inverse mapping from sparse observations $\boldsymbol{y}$ to the principal structure $\boldsymbol{m}$ using a deterministic autoencoder rather than a probabilistic model (e.g., a variational autoencoder). This choice is justified by the low uncertainty inherent in reconstructing $\boldsymbol{m}$ from $\boldsymbol{y}$ due to: (1) the strong sparse prior of $\boldsymbol{m}$ constraining the solution space since large-scale patterns vary smoothly and can be represented with few basis functions; (2) average pooling operation in the autoencoder promoting invariant and robust feature learning; and (3) our masked training strategy explicitly enhancing model’s robustness to noise and varying sensing patterns in $\boldsymbol{y}$. As a result, the posterior $p(\boldsymbol{m}|\boldsymbol{y})$ is highly concentrated, allowing us to approximate it as $p(\boldsymbol{m}|\boldsymbol{y})\approx\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$, where $\delta(\cdot)$ is a Dirac delta function and $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is the autoencoder output. This avoids instability from learning a full posterior variance, and still achieves accurate and robust reconstruction, as confirmed in our experiments.
2.2.1 The architecture of functional autoencoder
The encoder $\mathcal{E}$ maps a function $u:\Omega\rightarrow\mathbb{R}^{m}$ to a latent vector $z\in\mathbb{R}^{d_{z}}$. In developing the encoder–decoder architectures, we assume that $\mathcal{U}$ is a Banach space of functions that can be evaluated pointwise almost everywhere, with domain $\Omega\subset\mathbb{R}^{d}$ and codomain $\mathbb{R}^{m}$. In practice, we only have access to discretized representations $u(x)\in\mathcal{U}$, obtained by evaluating $u$ at a finite set of mesh points. Following the formulation in [li2020neural, li2020fourier], we parametrise the encoder as a kernel integral operator:
$\mathcal{E}(u;\zeta)=\rho\left(\int_{\Omega}\kappa\left(x,u(x);\zeta_{1}\right)\mathrm{d}x;\zeta_{2}\right)\in\mathcal{Z}=\mathbb{R}^{d_{z}},$ (2)
with $\kappa:\Omega\times\mathbb{R}^{m}\times\zeta_{1}\rightarrow\mathbb{R}^{l}$ parameterised as a neural network with hidden layers of width $l=64$, using GELU activation, and with $\rho:\mathbb{R}^{l}\times\zeta_{2}\rightarrow\mathbb{R}^{d_{z}}$ parametreised as the linear layer $\rho(v;\zeta_{2})=W^{\zeta_{2}}v+b^{\zeta_{2}}$, where $d_{z}$ denotes the dimension of latent vector $z$ and $\zeta=[\zeta_{1},\zeta_{2}]$ denotes the parameters of encoder. We augment $x\in\Omega$ with 16 random Fourier features to aid learning of high-frequency features [tancik2020fourier, li2020fourier]. After discretisation on data $\boldsymbol{u}=\{(x_{i},u(x_{i}))\}^{I}_{i=1}$, in which we approximate the integral over $\Omega$ by a normalised sum:
$\{(x_{i},u(x_{i}))|i=1,2,...I\}\rightarrow\rho(\text{pool}(\{\kappa(x_{i},u(x_{i});\zeta_{1})|i=1,2,...I\};\zeta_{2}))$ (3)
where pool is a pooling operation invariant to the order of its inputs—in our case, the average pooling. The average pooling operation reduces the discrepancy between latent representations derived from sparse and dense inputs, thereby endowing the functional autoencoder with the capability for accurate sparse-data reconstruction.
The decoder $\mathcal{D}$ maps a latent vector $z\in\mathcal{Z}$ back to a function $\mathcal{D}(z;\psi):\Omega\rightarrow\mathbb{R}^{m}$. To realize this mapping, we parameterize $\mathcal{D}$ with a coordinate-based neural network $\gamma:\mathcal{Z}\times\Omega\times\Psi\rightarrow\mathbb{R}^{m}$, whose hidden layers employ GELU activations. Accordingly, we have
$\mathcal{D}(z;\psi)(x)=\gamma(z,x;\psi)\in\mathbb{R}^{m}.$ (4)
In this construction, each spatial coordinate $x\in\Omega$ is enriched with 16 random Fourier features. Inspired by DeepONet [lu2021learning], the proposed architecture enables the decoded function $\mathcal{D}(z;\psi)$ to be evaluated on arbitrary meshes, with the computational cost scaling linearly in the number of discretization points.
2.2.2 Training objective and masked training
The central aim of training a functional autoencoder is to compress functional data $u$ into a low-dimensional latent code while retaining sufficient information to accurately reconstruct the original signal. Formally, the optimization problem is expressed as
[…]
2.3 Conditional diffusion model for detail generation
After reconstructing the principal structures with the functional autoencoder, a conditional diffusion model is trained to generate the small-scale components, defined as the difference between the ground truth and the reconstructed principal structures, conditioned on the principal structures. By defining the detail or residual $\boldsymbol{d}=\boldsymbol{u}-\boldsymbol{m}$, learning $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ is equivalent to learning $p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})$ as it is a deterministic relationship. Modeling the residual $\boldsymbol{d}$ simplifies the learning task, reduces complexity of modeling, and improves training stability and convergence. It also introduces an inductive bias, constraining the reconstructed $\boldsymbol{u}$ to remain close to $\boldsymbol{m}$, thereby restricting the solution space and mitigating ill-posedness. As discussed above, the functional autoencoder approximates $p(\boldsymbol{m}|\boldsymbol{y})$ with a Dirac delta function $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$. Consequently, the conditional diffusion model should be trained to model:
$p(\boldsymbol{u}|\boldsymbol{y})=\int p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}$ $=\int p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}\approx\int p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})\,\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))\,d\boldsymbol{m}=p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y}).$
2.3.1 Conditional denoising diffusion probabilistic model
Training of denoising diffusion probabilistic models (DDPMs) consists of a diffusion process and a denoising process. Diffusion process transforms $\boldsymbol{d}_{0}$ from the real data distribution $\boldsymbol{d}_{0}\sim p(\boldsymbol{d}_{0})$ into a pure Gaussian noise $\boldsymbol{d}_{t}\sim\mathcal{N}(0,\boldsymbol{I})$ by successively applying the following Markov diffusion kernel:
$p(\boldsymbol{d}_{t}|\boldsymbol{d}_{t-1})=\mathcal{N}\left(\boldsymbol{d}_{t};\sqrt{1-\beta_{t}}\,\boldsymbol{d}_{t-1},\,\beta_{t}\mathbf{I}\right),$ (6)
where $\{\beta_{t}\}^{T}_{t=1}$ is a pre-defined or learned noise variance schedule. We adopt a linear noise schedule where the variance $\beta_{t}$ increases linearly from $1\times 10^{-4}$ to $0.02$ over $T=1000$ diffusion steps. The marginal distribution at arbitrary timestep $t$ can be denoted as:
$p(\boldsymbol{d}_{t}|\boldsymbol{d}_{0})=\mathcal{N}\left(\boldsymbol{d}_{t};\sqrt{\bar{\alpha}_{t}}\,\boldsymbol{d}_{0},\,(1-\bar{\alpha}_{t})\,\mathbf{I}\right),$ (7)
where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s},\alpha_{t}=(1-\beta_{t})$. When $T\rightarrow\infty,\ p(\boldsymbol{d}_{t}|\boldsymbol{d}_{0})\approx\mathcal{N}(\mathbf{0},\mathbf{I})$. For sampling, $\boldsymbol{d}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{d}_{0}+(1-\bar{\alpha}_{t})\boldsymbol{\epsilon},\ \boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
Given $\boldsymbol{d}_{t}$, the denoising process aims to recover $\boldsymbol{d}_{0}$ by recursively learning the transition from $\boldsymbol{d}_{t-1}$ to $\boldsymbol{d}_{t}$, which is defined as the following Gaussian distribution:
$q_{\theta}(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t})=\mathcal{N}\left(\boldsymbol{d}_{t-1};\boldsymbol{\mu}_{\theta}(\boldsymbol{d}_{t},t),\boldsymbol{\Sigma}_{\theta}(\boldsymbol{d}_{t},t)\right),$ (8)
where parameters $\theta$ are optimized by a denoising network $\boldsymbol{\epsilon}_{\theta}$ that predicts $\boldsymbol{\mu}_{\theta}(\boldsymbol{d}_{t},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{\boldsymbol{d}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\,\boldsymbol{\epsilon}_{\theta}(\mathbf{\boldsymbol{d}}_{t},t)\right)$. As for $\boldsymbol{\Sigma}_{\theta}(\boldsymbol{d}_{t},t)$, we use the closed-form posterior variance $\sigma_{t}^{2}=\beta_{t}\frac{1-\bar{\alpha}_{t-1}}{1-\bar{\alpha}_{t}}$ as proposed in the original denoising diffusion probabilistic model formulation [ho2020denoising], instead of predicting it via neural network.
However, this unconditional scheme fails to preserve the uniqueness of reconstruction due to the lack of guidance during generation. Therefore, $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is injected as the condition for the denoising process:
$q_{\theta}(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))=\mathcal{N}(\boldsymbol{d}_{t-1};\boldsymbol{\mu}_{\theta}(\boldsymbol{d}_{t},t,\hat{\boldsymbol{m}}(\boldsymbol{y})),\sigma^{2}_{t}\mathbf{I}),$ (9)
where $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is injected by concatenating with the input $\boldsymbol{d}_{t}$ along the channel dimension.
To generate samples from the trained model, the reverse process is iteratively performed starting from a Gaussian noise $\boldsymbol{d}_{T}\sim\mathcal{N}(0,\mathbf{I})$ as follows:
$\mathbf{\boldsymbol{d}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{\boldsymbol{d}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{\boldsymbol{d}}_{t},t,\hat{\boldsymbol{m}}(\boldsymbol{y}))\right)+\sigma_{t}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (10)
This process is repeated until $t=1$, at which point $\boldsymbol{d}_{0}$ is taken as the final reconstructed sample.
2.3.2 Manifold constrained gradient
The conditional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by training the conditional diffusion model. The $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ should be modeled to further preserve the uniqueness of reconstruction and impose measurement consistency. However, directly model the condtional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ is impossible due to the immense variability of sparse measurement patterns. We model it during inference instead of training of conditional diffusion model. Specifically, the Manifold Constrained Gradient (MCG) [chung2022improving] is used to inject an additional gradient term to guide the generation towards observed measurements without retraining the model. This gradient refers to the score function in score-based generative models (SGMs) [song2020score], defined as the gradient of the log-density of the data distribution $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
Although DDPMs and SGMs were originally proposed under different formulations—predicting added noise in DDPMs and estimating score functions in SGMs—it has been shown that they are theoretically equivalent under certain parameterizations [song2020score]. Specifically, when a DDPM is trained to predict the Gaussian noise $\boldsymbol{\epsilon}_{\theta}$ added at each step, the network implicitly learns the score function of the perturbed data distribution $p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ up to a scaling factor:
$\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))\approx-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{d}_{t},t,\hat{\boldsymbol{m}}(\boldsymbol{y})),$ (11)
Therefore, the sampling equation in Eq. 10 can be reformulated as:
$\mathbf{\boldsymbol{d}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{\boldsymbol{d}}_{t}+(1-\alpha_{t})\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))\right)+\sigma_{t}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (12)
This connection enables MCG to be directly applied to DDPMs. Specifically, the Bayes rule $p(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})=\frac{p(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))p(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))}{p(\boldsymbol{y}|\hat{\boldsymbol{m}}(\boldsymbol{y}))}$ leads to:
$\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})=\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))+\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y})),$ (13)
where $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by conditional diffusion model. To approximate the likelihood term $p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$, we assume that it follows a Gaussian form, i.e.,
$p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))\propto\exp\left(-\frac{1}{2\sigma_{c}^{2}}\|\boldsymbol{y}-\mathcal{M}(\hat{\boldsymbol{m}}(\boldsymbol{y})+\boldsymbol{d}_{t})\|_{2}^{2}\right),$ (14)
where $\mathcal{M}(\cdot)$ denotes the mask operator and $\sigma_{c}^{2}$ denotes the variance associated with measurement noise. $\sigma_{c}^{2}$ is set as 10000 in all experiments of this study. Under this assumption, its gradient with respect to $\boldsymbol{d}_{t}$ yields:
[…]
Therefore, the sampling formula without retraining the model is given as:
[…]
However, it drives the intermediate states at each denoising step to match the observations, which pulls the sample path away from the learned data manifold. Chung et al. [chung2022improving] reveal that the Bayes optimal denoising step from the Tweedie’s formula [robbins1992empirical] leads to a preferred condition for diffusion model both empirically and theoretically. As a result, Tweedie’s formula is introduced here to impose the measurement consistency while improving the generation accuracy. In the case of Gaussian noise, a classic result of Tweedie’s formula tells us that one can achieve the denoised result by computing the posterior expectation:
[…]
where the noise is modeled by $\tilde{\boldsymbol{d}}\sim\mathcal{N}(\boldsymbol{d},\sigma^{2}\mathbf{I})$. Considering the conditional diffusion model where the forward step is modeled as $\boldsymbol{d}_{t}\sim\mathcal{N}\left(\sqrt{\bar{\alpha}_{t}}\,\boldsymbol{d}_{0},\,(1-\bar{\alpha}_{t})\,\mathbf{I}\right)$, the Tweedie’s formula can be rewritten as:
$\mathbb{E}[\boldsymbol{d}_{0}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y})]=\frac{1}{\sqrt{\bar{\alpha}_{t}}}\big(\boldsymbol{d}_{t}+(1-\bar{\alpha}_{t})\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))\big)$ (17)
By replacing $\boldsymbol{d}_{t}$ in Eq.(15) with the posterior expectation computed by Tweedie’s formula $\hat{\boldsymbol{d}}_{0}(\boldsymbol{d}_{t})=\mathbb{E}[\boldsymbol{d}_{0}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y})]$, the sampling formula for the trained conditional diffusion model is given as:
[…]
2.3.3 Mask-cascade training
We propose a mask-cascade training strategy to improve the robustness of the conditional diffusion model under varying sparsity patterns. A key challenge is that the functional autoencoder, when provided with different sparse samplings of the same field, reconstructs slightly different principal structures. Since it is infeasible to enumerate all possible reconstructions for training, directly using a fixed dataset would restrict the generalization ability of the diffusion model.
To address this issue, our strategy explicitly integrates the stochasticity of sparse sampling into training the conditional diffusion model. As illustrated in Fig. 1(b), once the functional autoencoder is fully trained and its parameters are frozen, the training of the conditional diffusion model proceeds as follows: in each training step, we randomly generate a sparse mask and apply it to the ground-truth field $\boldsymbol{u}$ to obtain a partial observation $\boldsymbol{y}$. This observation is then passed through the functional autoencoder to reconstruct an approximate principal structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$. The conditional diffusion model is then conditioned on $\hat{\boldsymbol{m}}(\boldsymbol{y})$, and trained to generate the fine-scale component $\boldsymbol{d}=\boldsymbol{u}-\hat{\boldsymbol{m}}(\boldsymbol{y})$. By repeatedly sampling different masks across training steps, the conditional diffusion model is continuously exposed to a diverse ensemble of approximate principal structure reconstructions. This effectively augments the training distribution of conditioning inputs without the need for explicitly precomputing or storing all possible variants. As a result, the conditional diffusion model learns to model the conditional distribution $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$, thereby achieving improved robustness and consistency across varying sparsity patterns.
As shown in Fig.1(c), the inference process of autoencoder-diffusion cascade proceeds as follows. Given arbitrary sparse observations $\boldsymbol{y}$, the functional autoencoder first reconstructs the principal structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$. This reconstructed structure is then provided as a condition to the conditional diffusion model, which generates the fine-scale details $\hat{\boldsymbol{d}}$ through the manifold constrained gradient-enhanced sampling scheme in Eq. (18). Finally, the full-field reconstruction is obtained by combining the two components $\hat{\boldsymbol{u}}=\hat{\boldsymbol{m}}(\boldsymbol{y})+\hat{\boldsymbol{d}}$. It is worth noting that, during inference, Cas-Sensing can still achieve reliable reconstruction even when the sampling rate of sparse observations differs significantly from the rates used in mask-cascade training.
(a) The masked training scheme of functional autoencoder. The trained functional autoencoder is to approximate $p(\boldsymbol{m}|\boldsymbol{y})$ with $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$ as it learns the deterministic mapping between $\boldsymbol{y}$ and $\boldsymbol{m}$. $x_{e}$ and $u_{e}$ denote the coordinates and field values fed into the encoder, while $x_{d}$ and $u_{d}$ represent the coordinates and reconstructed values produced by the decoder.
(b) The mask-cascade training scheme of conditional diffusion model, which ensure the conditional diffusion model’s accurate capturing $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
(c) The inference scheme of Cas-Sensing based on an autoencoder-diffsusion cascade, which is the process of sampling from $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$. Manifold Constrained Gradient (MCG) is used to impose measurement consistency and sample from $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ without retraining the model.
Figure 1: Overview of the proposed Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade.
3 Results
3.1 Reconstructing cylinder flow velocity fields with simulation data
To evaluate the effectiveness of the proposed Cas-Sensing based on an autoencoder-diffusion cascade in reconstructing full-field data from randomly sparse observations under varying geometric boundary conditions, we first constructed a 2D cylinder flow dataset. This dataset captures incompressible fluid dynamics around a 2D circular cylinder within a channel:
$\nabla\cdot\mathbf{u}$ $=0,$ (19)
$\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}$ $=\mu\nabla^{2}\mathbf{u}-\frac{1}{\rho}\nabla p.$ (20)
Specifically, the computational domain is defined as a rectangular region of size $1.6\times 0.6$, within which both the cylinder center and radius are randomly sampled. This procedure yields a dataset comprising 96 distinct boundary configurations, each associated with 100 temporal snapshots. All flow fields are interpolated onto a uniform grid of size $256\times 96$. The visualization of a sample from the Cylinder Flow is presented in Fig. 2. For the purpose of model training and evaluation, the dataset is partitioned based on boundary configurations rather than temporal snapshots, ensuring that the test set contains unseen geometrical variations. Specifically, 76 boundary configurations, corresponding to a total of 7600 flow field snapshots, are employed as the training set, while the remaining 20 configurations, consisting of 2000 snapshots, are reserved exclusively for testing.
Figure 2: An example of 2D cylinder flow velocity fields. The dataset contains 96 different boundary configurations of circular cylinders, each associated with 100 temporal snapshots.
We first employ a functional autoencoder to learn compact representations of the circular cylinder flow fields. To encourage robustness, the model is trained in a masked manner, where only $50\%$ of the grid points in each snapshot are provided to the encoder and the task is to reconstruct the full field from this partial information. To investigate the latent representation learned by the functional autoencoder, we first visualize the extracted latent vectors of the cylinder flow dataset. Specifically, the 32-dimensional latent vectors are projected into a two-dimensional space using the t-distributed Stochastic Neighbor Embedding (t-SNE) technique [maaten2008visualizing]. Each point in Fig. 3 corresponds to the latent variable of a single flow snapshot. It can be observed that the training data associated with different internal boundary configurations are well separated in the latent space, and the latent vectors of snapshots from the same configuration form circular patterns. This circular structure reflects the inherent periodicity of the velocity fields in cylinder wake flows. Moreover, the latent vectors corresponding to the test dataset exhibit the same characteristic structures, indicating that the functional autoencoder has effectively captured both the temporal and geometric features of the flow fields.
Figure 3: t-SNE visualization of latent representations learned by the functional autoencoder for cylinder flow fileds.
To further assess the model’s reconstruction capability under varying levels of data sparsity, we select a representative flow field from the test set and apply 100 randomly generated masks with sampling ratios of $50\%,10\%,3\%,1\%$, and $0.5\%$. The masked inputs are then mapped through the trained functional autoencoder to recover the complete field on the $256\times 96$ grid. The distributions of the resulting reconstruction errors, quantified by the root mean square error (RMSE), are illustrated in Fig. 4. Quantitative results confirm the strong robustness of the functional autoencoder in reconstructing cylinder wake flows under varying levels of input sparsity. Even when the available input points are reduced from $50\%$ to $10\%$, the mean RMSE of the reconstructions remains nearly constant (around 0.291), with only a negligible change in variance. Noticeable degradation occurs only when the input ratio drops below $3\%$, where the RMSE increases modestly (to about 0.31 at $0.5\%$ input). These results demonstrate that the autoencoder is capable of faithfully recovering the dominant flow features even with extremely sparse measurements, underscoring its resilience to data sparsity and strong generalization ability.
Figure 4: The reconstructed principal structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
After training the functional autoencoder, we freeze its parameters and employ it to provide structural conditions for a conditional diffusion model. In mask-cascade traing of the conditional diffusion model, only $0.5\%$ of the grid points are supplied as input to the functional autoencoder, which then reconstructs the principal flow structures serving as guidance for the conditional diffusion model. During testing, these reconstructed structures are combined with the sparse measurements to enable the conditional diffusion model to recover the complete flow fields. Representative results, shown in Fig. 5(a), indicate that the generated fields reproduce the ground truth with high accuracy, capturing both the periodic vortex shedding and the influence of varying boundary configurations.
To further probe reconstruction reliability, we select a test snapshot and impose random masks with different sampling ratios ($0.1\%,0.5\%,1\%,3\%$, and $5\%$). For each ratio, 100 independent reconstructions are carried out, and the kernel density distributions of RMSE are computed, as presented in Fig. 5(b). The results demonstrate a clear trend: at extremely sparse ratios such as $0.1\%$, the reconstructions remain accurate but exhibit slightly higher variance; as the ratio increases to $1\%$, both the mean error and uncertainty decrease significantly, reflecting enhanced stability of the generated fields. Beyond $3\%$, the performance saturates, with the error approaching a lower bound and the variance remaining minimal. Importantly, although the model is trained using only $0.5\%$ observations, it generalizes effectively to both sparser and denser cases, consistently delivering faithful reconstructions. This robustness is attributed to the randomization inherent in the mask-cascade training, which exposes the model to diverse observation patterns, and to the incorporation of manifold constraints, which enforce consistency between generated flows and sparse inputs.
(a) The generated full-field reconstructions with $0.5\%$ input point ratio. Each reconstruction is a sample from $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$.
(b) The mean full-field reconstructions of 100 generations with different input point ratios and corresponding kernel density estimates. The mask is fixed at each input point ratio.
Figure 5: The full-field reconstruction results of cylinder flow fields.
3.2 Reconstructing sea surface wave height fields with stereo data
In this subsection, we demonstrate the Cas-Sensing on stereo image data of sea surface wave height fields for reconstruction from sparse measurements, highlighting its utility in complex real-world engineering. Stereo imaging measurement of the sea surface height is based on single snapshots or time records captured by a pair of synchronized and calibrated cameras. The example of stereo-image pair process by Wave Acquisition Stereo System (WASS) is presented in Fig. 6. The data set used in this study is Acqua Alta stereo data set including 8000 image frames, recording a half hour sea evaluation with an imaging system installed on the north-east side of the Acqua Alta oceanographic research tower.
Figure 6: An example of stereo-image pair (from the Acqua Alta tower in the northern Adriatic Sea, Italy) and 3-D wave field on top of the right image (the colorscale is proportional to the sea surface height)[guimaraes2020data].
First, we train a functional autoencoder to reconstruct the principal structures of sea surface wave fields using 7000 image frames, with the remaining 1000 frames held out as a test set. During masked training, $50\%$ of the data points in each image are provided to the encoder as input for feature extraction, while the model is trained to reconstruct the complementary $50\%$ of the points. To evaluate the reconstruction performance, we investigate the sensitivity of the model to different input point ratios. Specifically, we fix an arbitrary sample from the held-out set and generate 100 distinct masks with point ratios of $50\%,10\%,3\%,1\%$, and $0.5\%$. Each masked input is then encoded and decoded on the full grid, and kernel density estimates of the reconstruction root mean square error (RMSE) are computed, as presented in Fig. 7.
From Fig. 7, it is evident that when the input ratio is above approximately $3\%$, the functional autoencoder achieves highly accurate and stable reconstructions of the principal structures, with RMSE means remaining close to 0.40 and variances below 0.015. As the input ratio decreases further, the reconstruction quality gradually degrades. At $1\%$ input, for example, the mean RMSE increases to about 0.43 and the variance nearly doubles, and at $0.5\%$ input the mean RMSE reaches approximately 0.45 with the largest spread in the error distribution. Despite this deterioration, the reconstructions still preserve the correct large-scale patterns of the wave field, highlighting the strong robustness of the functional autoencoder to severe data sparsity in capturing the dominant structures.
Figure 7: The reconstructed principal structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
When the functional autoencoder is trained and fixed, it is used to output principal structures according to sparse measurements as conditions to train a conditional diffusion model with the proposed mask-cascade training strategy. The input point ratio of the fixed functional autoencoder is set as $0.5\%$ to enhance the robustness of conditional diffusion model to sparsity. To test the generation results, given $0.5\%$ randomly selected observation points from each test sample, the functional autoencoder reconstructs the corresponding principal structures. These structures are then used as conditions to input the conditional diffusion model while manifold constrained gradients are incorporated into the sampling process to ensure consistency between the generated full-field data and the observed $0.5\%$ points. As shown in Fig. 8(a), the generated fields exhibit high fidelity to the ground truth. It is important to note that each generated realization corresponds to a single sample from $p(\boldsymbol{u}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$. To quantitatively evaluate the uncertainty of the generative model, we further fix a test sample and generate random masks at different observation ratios ($0.2\%,0.5\%,1\%,3\%$, and $5\%$). After fixing these masks, we perform 100 independent generations, and compute the kernel density estimates of the resulting RMSE distributions. The results, shown in Fig. 8, where the mean field of generations match the ground truth very well. As the percentage of observed points increases from $0.2\%$ to $1\%$, the mean RMSE decreases notably, indicating improved reconstruction quality with denser observations. In particular, the drop in both mean error and variance around the $1\%$ input level suggests that even a relatively small increase in observations can significantly stabilize the generation. When the ratio is further increased to $3\%$ and $5\%$, the mean RMSE continues to decrease slightly, approaching a lower bound, while the variance remains at a low level since the principal structures reconstruction results are more stable under these ratios, reflecting both accuracy and robustness. It is worth noting that although the conditional diffusion model is trained with $0.5\%$ input points, it generalizes well to different input ratios and consistently delivers satisfactory reconstructions even with only $0.2\%$ points. This can be attributed to two factors. First, during mask-cascade training, the conditions are generated by random sampling, which enhances the robustness of the conditional diffusion model to diverse and sparse conditions. Second, the manifold-constrained gradient further enforces consistency between the generated fields and the sparse observations while keeping the sampling trajectory on the data manifold, which ensures reliable reconstruction of the physical fields even under varying levels of sparsity.
(a) The generated full-field reconstructions with $0.5\%$ input point ratio. Each reconstruction is a sample from $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$.
(b) The mean full-field reconstructions of 100 generations with different input point ratios and corresponding kernel density estimates. The mask is fixed at each input point ratio.
Figure 8: The full-field reconstruction results of sea surface wave fields.
3.3 Reconstructing global sea surface temperature fields with reanalysis data
We also employed the daily global Sea Surface Temperature (SST) reanalysis dataset provided by the Copernicus Marine Service [SST_dataset] to evaluate the reconstruction performance of the proposed Cascaded Sensing method with sparse observations. We selected data spanning three years, from June 1, 2022, to May 30, 2025. Due to the high spatial resolution of the original dataset (4320 × 2040), the data were first downsampled to 1440 × 640. The downsampled fields were then evenly partitioned into 15 subregions, each with a resolution of 480 × 128, as illustrated in Fig. 9, forming SST dataset. For model training and evaluation, the first $80\%$ of the data were used as the training set, and the remaining $20\%$ were reserved for testing.
Figure 9: The partition of global sea surface temperature field data set.
We begin by training a functional autoencoder on the dataset to learn a compact representation of the global sea surface temperature field. During the masked training, $50\%$ of the data points in each frame are randomly selected as inputs to the encoder, while the remaining $50\%$ are used as targets for the decoder. As illustrated in Fig. 10(a) (Condition), the trained functional autoencoder is capable of accurately reconstructing the large-scale background patterns from only partial observations, capturing the essential spatial variability of the ocean surface temperature fields. Once the autoencoder is fully trained, it is frozen and integrated into the Cascaded Sensing framework to provide real-time background field reconstructions that serve as conditioning information for the conditional diffusion model during mask-cascade training. Specifically, a sparse subset of points ($0.5\%$) is randomly sampled at each step during training, from which the functional autoencoder reconstructs a smooth background field. The conditional diffusion model then learns to refine this background and recover the full field. After 800 epochs of training, the conditional diffusion model is well-trained to reconstruct fine-scale details based on conditions and extremely sparse measurements.
To evaluate the model’s generalization ability under more challenging observation conditions, we randomly sample only $0.1\%$ of the data points from the 15 predefined global subregions, a sampling ratio that is five times lower than that used during training. A representative reconstruction (RMSE = 0.0144) is shown in Fig. 10(a). Remarkably, as evident in the zoomed-in views, the Cascaded Sensing framework not only reconstructs the global background temperature field with high accuracy but also successfully captures fine-scale turbulent structures, which are typically difficult to recover from such sparse observations.
To systematically assess the robustness of the proposed approach, we further conduct experiments under a range of observation ratios ($5\%,3\%,1\%,0.5\%,0.1\%$). For each ratio, sparse points are randomly selected and fixed, and the model generates 100 samples to characterize the reconstruction variability. The resulting mean and standard deviation of RMSE are summarized in Fig. 10(b). As expected, both the mean and standard deviation of RMSE increase as the observation ratio decreases. Nevertheless, the errors remain consistently low across all tested ratios, underscoring the robustness, stability, and generalization capability of the cascaded sensing framework in extremely sparse sensing scenarios.
(a) The generated full-field reconstructions with $0.1\%$ input point ratio. Each reconstruction is a sample from $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$.
(b) The mean full-field reconstructions of 100 generations with different input point ratios and corresponding RMSE. The mask is fixed at each input point ratio.
Figure 10: The full-field reconstruction results of global sea temperature fields.
4 Discussion
In this study, we conduct a comprehensive evaluation of the proposed Cascaded Sensing (Cas-Sensing) framework under a broad range of scenarios characterized by different sensor configurations, sampling sparsities, and geometric boundaries. Our results clearly show that the hierarchical structure of Cas-Sensing can substantially mitigate the ill-posedness of reconstructing multi-scale physical fields from extremely sparse measurements. By decomposing the reconstruction process into two stages—large-scale structure recovery via a functional autoencoder and fine-scale detail generation through a conditional diffusion model—the proposed method achieves accurate and stable reconstructions with high data fidelity.
A central strength of this cascaded design lies in its ability to reduce the reconstruction ambiguity caused by sparse and irregular inputs. By first establishing reliable coarse-scale conditioning, the diffusion model no longer needs to infer the entire field from incomplete observations, but focuses on capturing the missing small-scale structures. This hierarchical formulation enhances reconstruction stability and reduces the variability across different sampling patterns. Moreover, the mask-cascade training strategy significantly improves the model’s robustness and generalization, enabling it to adapt to diverse sensor layouts without the need for retraining.
These advantages make Cas-Sensing particularly effective for scientific applications. For example, in fluid dynamics, accurate flow field reconstruction from sparse sensors is crucial for monitoring and control; in climate and ocean science, reconstructing global or regional temperature fields from sparse measurements enables more reliable environmental assessment; and in geophysical applications, the ability to generalize across sampling configurations supports practical deployment in complex field settings.
Overall, Cas-Sensing provides a flexible and efficient framework for reconstructing multi-scale fields. By leveraging hierarchical representation learning and generative modeling, it delivers both accuracy and robustness across a variety of reconstruction scenarios, offering a promising pathway for advancing data-driven scientific sensing and prediction.
5 Conclusion and future work
In this study, we have presented Cascade Sensing (Cas-Sensing) as an effective framework for reconstructing full-field data from highly sparse observations. By leveraging an autoencoder-diffusion cascade framework, Cas-Sensing is capable of producing accurate and coherent reconstructions even when the input measurements are extremely limited. The mask-cascade training strategy, combined with manifold constrained gradient, enhances the robustness of the model to diverse sparse conditions and ensures the consistency of the reconstructed physical fields. These results demonstrate the potential of Cas-Sensing as a powerful tool for efficient, data-driven sensing and predictive modeling in complex physical systems.
Nevertheless, several limitations remain in the current implementation. Firstly, the present study focuses exclusively on two-dimensional spatial domains, leaving the extension to higher-dimensional settings unexplored. Scaling Cas-Sensing to three-dimensional or higher-dimensional problems may introduce additional computational and modeling challenges that warrant careful investigation. Secondly, the current framework addresses only spatial reconstruction, without explicitly modeling temporal dynamics. Consequently, the method cannot fully capture spatiotemporal evolution in dynamic systems, which may limit its applicability in real-time monitoring or predictive scenarios. Thirdly, the generative component currently relies on a vanilla DDPM, which, while effective, may impose constraints on sampling speed and computational efficiency. Alternative generative models, such as flow matching, could potentially accelerate sampling and further improve reconstruction fidelity.
Future research should aim to overcome these limitations. Extending Cas-Sensing to higher-dimensional domains would broaden its applicability to more complex physical systems, including 3D flow fields or volumetric sensor networks. Integrating temporal modeling within the cascade framework could enable spatiotemporal reconstruction, capturing dynamic evolution in addition to spatial correlations, and allowing the framework to support real-time prediction tasks. Moreover, the adoption of more advanced generative models could enhance sampling efficiency without sacrificing accuracy, making Cas-Sensing more practical for large-scale deployments. Collectively, these developments would not only enhance the versatility and scalability of Cas-Sensing but also reinforce its potential as a robust and reliable tool for data-driven sensing, reconstruction, and prediction in complex physical environments.

### v2

@@
 Abstract
-Reconstructing full fields from extremely sparse and random measurements is a longstanding ill-posed inverse problem.
-A powerful framework for addressing such challenges is hierarchical probabilistic modeling, where uncertainty is represented by intermediate variables and resolved through marginalization during inference.
-Inspired by this principle, we propose Cascaded Sensing (Cas-Sensing), a hierarchical reconstruction framework that integrates an autoencoder–diffusion cascade.
-First, a neural operator-based functional autoencoder reconstructs the dominant structures of the original field—including large-scale components and geometric boundaries—from arbitrary sparse inputs, serving as an intermediate variable.
-Then, a conditional diffusion model, trained with a mask-cascade strategy, generates fine-scale details conditioned on these large-scale structures.
-To further enhance fidelity, measurement consistency is enforced via the manifold constrained gradient based on Bayesian posterior sampling during the generation process.
-This cascaded pipeline substantially alleviates ill-posedness, delivering accurate and robust reconstructions.
-Experiments on both simulation and real-world datasets demonstrate that Cas-Sensing generalizes well across varying sensor configurations and geometric boundaries, making it a promising tool for practical deployment in scientific and engineering applications.
+Reconstructing full fields from extremely sparse and random measurements constitutes a fundamentally ill-posed inverse problem, in which deterministic end-to-end mappings often break down due to intrinsic non-uniqueness and uncertainty.
+Rather than treating sparse reconstruction as a regression task, we recast it as a hierarchical probabilistic inference problem, where uncertainty is explicitly represented, structured, and progressively resolved.
+From this perspective, we propose Cascaded Sensing (Cas-Sensing) as a general reconstruction paradigm for multi-scale physical fields under extreme data sparsity.
+Central to this paradigm is the introduction of an explicit intermediate representation that decomposes the original ill-posed problem into two substantially better-conditioned subproblems.
+First, a lightweight neural-operator-based functional autoencoder infers a coarse-scale approximation of the target field from sparse observations acting as an explicit intermediate variable.
+Rather than modeling multiple scales jointly, this intermediate estimate is deterministically fixed and subsequently used as the sole conditioning input to a conditional diffusion model that generates refined-scale details, yielding a cascaded inference structure with clearly separated reconstruction responsibilities.
+To ensure robustness under diverse sensing patterns, the diffusion model is trained using a mask-cascade strategy, which exposes it to a distribution of imperfect conditioning structures induced by extreme sparsity.
+During inference, measurement consistency is enforced through manifold-constrained gradients within a Bayesian posterior framework, ensuring fidelity to sparse observations while preserving data manifold coherence.
+This cascaded probabilistic formulation substantially alleviates ill-posedness, enabling accurate and stable reconstructions even under extreme sparsity.
+Extensive experiments on both simulated and real-world datasets demonstrate that Cas-Sensing generalizes effectively across sensor layouts and sparsity levels.
+These results suggest that Cas-Sensing provides not only an effective reconstruction tool, but also a principled and extensible framework for data-driven scientific sensing and inference in complex physical systems.
 keywords:
-Field reconstruction; Probabilistic modeling; Cascade pipeline; Functional autoencoder; Diffusion model.
+Full-Field reconstruction; Probabilistic modeling; Cascade pipeline; Functional autoencoder; Diffusion model.
 1 Introduction
-Despite advances in sensor technology enabling large-scale deployments in physical systems, the number of monitoring sites is often constrained by high setup and maintenance costs.
+Despite advances in sensor technology enabling coarse-scale deployments in physical systems, the number of monitoring sites is often constrained by the high cost of setup and maintenance.
 Therefore, spatial field reconstruction from limited local sensor information is a major challenge of complex physical systems, such as astrophysics [akiyama2019first], geophysics [kondrashov2006spatio, carrassi2018data, manohar2018data], atmospheric science [alonso2010novel, mishra2014compressed], and fluid dynamics [vinuesa2023transformative, buzzicotti2023data].
-Due to the limited sensors, the reconstruction of multi-scale full-field data from extremely sparse and random observations is inherently an ill-posed problem, where the results often suffers from issues related to existence, uniqueness, and stability [hadamard2014lectures, breiding2023algebraic].
-Compressed sensing has been widely applied to address this challenge.
-It relies on the assumption that signals are sparse in a specific transform domain, and reconstructs full-field data from undersampled observations by solving optimization problems [candes2006robust].
+Due to the deployment of limited sensors, the reconstruction of multi-scale full-field data from extremely sparse and random observations is inherently an ill-posed problem with high uncertainty [hadamard2014lectures, breiding2023algebraic].
+Compressed sensing has been widely applied to address this type of problems.
+It relies on the assumption that signals are sparse in a specific transform domain such that uncertainty is suppressed, and reconstructs full-field data from undersampled observations by solving optimization problems [candes2006robust].
 To enhance reconstruction performance, compressed sensing variants combined with deep learning have been proposed, aiming to automatically learn effective sparse representations and reconstruction mappings in a data-driven manner [sun2017intelligent, shi2019image, ni2020deep, xu2022sparse].
@@
 However, such mappings are typically trained with a specific reorganization form of sparse data (for example, patch or voronoi tessellation) and geometric boundary.
-When encountering unseen sparse measurements—such as data acquired from a different sensor configuration—or when facing new geometric boundaries, these models often fail to produce accurate reconstructions and require retraining to adapt to the new scenarios.
+When encountering unseen sparse measurements—such as data acquired from a different sensor deployment configuration, or when facing new geometric boundaries, these models often fail to produce accurate reconstructions and require retraining to adapt to the new scenarios.
 As a result, the applicability of these methods is generally limited to controlled or narrowly defined settings.
 They often struggle with generalization when applied to complex, real-world environments.
-Moreover, when the available observations become extremely sparse, the inherent ill-posed nature of the reconstruction problem severely impedes the ability to establish a stable and reliable mapping from sparse inputs to the full-field data [bell1978solutions, candes2011tight, arridge2019solving].
+From an inverse problem perspective, the limitation of end-to-end mappings is fundamentally rooted in the intrinsic uncertainty of sparse reconstruction.
+Under severely incomplete observations, the mapping from sparse measurements to full-field solutions is generally non-unique, meaning that multiple physically plausible fields may correspond to the same set of measurements.
+This ambiguity is a defining characteristic of ill-posed inverse problems and cannot be resolved by deterministic regression alone [bell1978solutions, arridge2019solving].
+As a result, learning a single deterministic mapping inevitably collapses the underlying solution space, often leading to instability, overfitting to specific sensing configurations, and poor generalization when observations become sparser or deviate from the training distribution.
+In contrast, probabilistic generative modeling provides a natural mechanism for representing and sampling from the conditional distribution of feasible solutions, rather than committing to a single point estimate.
+By modeling the data generation process itself, generative models are able to capture uncertainty and variability arising from incomplete observations, making them particularly suitable for highly underdetermined reconstruction tasks.
 Generative models, such as generative adversarial networks [goodfellow2020generative] and diffusion models [ho2020denoising, song2020score], have shown remarkable capabilities in producing high-quality, diverse data, with broad applications in vision [wang2023dr2, miao2024waveface, zhao2023towards], biomedicine [cao2024high, yu2020medical, nguyen2024sequence], and scientific computing [lienen2306zero, ruhling2023dyffusion, li2024synthetic].
@@
 One line of work learns supervised conditional distributions between full-field low-resolution and high-resolution data pairs [deng2019super, shu2023physics].
-However, due to the immense variability of sparse measurement patterns, directly modeling the distribution of full-field data conditioned on sparse observations is practically infeasible, as it is impossible to cover all sparse-to-full field mappings during training.
+However, due to the immense variability of sparse measurement patterns, directly modeling the distribution of full-field data conditioned on sparse observations is practically infeasible, as it is impossible to numerate all sparse-to-full field mappings during training.
 Another approach adopts a reverse diffusion step followed by a projection-based measurement consistency step [choi2021ilvr, chung2022come, kadkhodaie2021stochastic] or a Bayesian sampling step [du2024conditional, li2024learning] after pretraining an unconditional diffusion model to ensure the generated samples match sparse observations.
@@
 Moreover, it extends diffusion models to function spaces by coupling a function autoencoder with latent diffusion, enabling physics-informed generative modeling across different discretizations.
-While powerful for capturing global structures, this latent-space generation strategy introduces an information bottleneck: fine-scale details may be lost during encoding and cannot be fully recovered during decoding, which can limit its effectiveness in detail reconstruction from sparse observations.
-Cascading pipeline of generative model is an effective way to improve the generation quality in case without strong conditioning information [nichol2021improved, saharia2022image, ho2022cascaded].
+While this latent-space formulation is effective for capturing global structures, the unified generation of multi-scale components within a single latent representation constrains the model’s ability to independently regulate fine-scale details, particularly under extremely sparse conditioning.
+A cascade pipeline of generative models is an effective way to improve the generation quality in cases without strong conditioning information [nichol2021improved, saharia2022image, ho2022cascaded].
 In computer vision, this approach involves training a sequence of models at progressively higher resolutions: a base model generates low-resolution samples, which are then refined by super-resolution models [ho2022cascaded].
@@
 Moreover, individual models can be optimized independently, with architecture choices tailored to each resolution for optimal pipeline performance.
-Inspired by the cascading pipeline, the Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade is proposed to achieve accurate and robust multi-scale fields reconstruction from extremely sparse measurements.
-In Cas-Sensing, by viewing the full-field data as functional data, an neural operator-based functional autoencoder is first trained [bunker2024autoencoders] to extract the latent features of data, thus having the ability of robustly recovering the principal structures (large-scale components with geometric boundaries) from arbitrary sparse measurements.
-Then a conditional diffusion model is trained with the conditions—the reconstructed principal structures—to only focus on generating the small-scale details which are the difference between ground truth and conditions.
+As shown in Fig.
+1, inspired by the hierarchical probabilistic decomposition via marginalization, the Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade is proposed to achieve accurate and robust multi-scale fields reconstruction from extremely sparse measurements.
+In Cas-Sensing, by viewing the full-field data as functional data (data sampled from continuous functions), a neural operator-based functional autoencoder is first trained [bunker2024autoencoders] to extract the latent features of data, thus having the ability of robustly recovering the coarse-scale structures from arbitrary sparse measurements.
+Then a conditional diffusion model is trained with the conditions—the reconstructed coarse-scale structures—to only focus on generating the fine-scale details, which are the residuals between ground truth and conditions.
 To enhance the robustness of diffusion model, we introduce a Mask-Cascade Training (MCT) approach, which applies different random masks during training to generate diverse conditions from a fixed ground-truth field.
 After training the conditional diffusion model, we employ the Manifold Constrained Gradient (MCG) [chung2022improving] during sampling to ensure that the generated full-field data not only matches the sparse observations at measurement locations, but also remains consistent with the underlying data manifold.
-The proposed cascade pipeline effectively alleviates the ill-posedness of the original reconstruction problem by progressively recovering multi-scale components in a staged manner, thereby enabling Cas-Sensing to achieve high accuracy, strong generalizability, and enhanced robustness in reconstruction tasks.
+The proposed cascade pipeline effectively alleviates the ill-posedness of the reconstruction problem by progressively recovering multi-scale components in a staged manner, thereby enabling Cas-Sensing to achieve high accuracy, strong generalizability, and enhanced robustness in reconstruction tasks.
 The remainder of the paper is organized as follows.
@@
 2.1 Problem formulation
+Figure 1: The overview of Cascaded Sensing, where $\boldsymbol{d}=\boldsymbol{u}-\boldsymbol{m}$ and MCG denotes Manifold Constrained Gradient for imposing the measurement consistency.
 The problem of recovering the multi-scale full field $\boldsymbol{u}\in\mathbb{R}^{m}$ from a vector of sparse sensor measurements $\boldsymbol{y}\in\mathbb{R}^{n}$ can be formulated as modeling the conditional distribution $p(\boldsymbol{u}|\boldsymbol{y})$.
 However, when the observations are excessively sparse, directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$ is ill-posed, as the solution becomes highly non-unique.
-To address this challenge, we propose Cas-Sensing, a cascaded framework that mitigates the ill-posedness by leveraging the following marginalization rule:
+To address this challenge, we introduce an intermediate variable $\boldsymbol{m}$ into this probabilistic framework, thus forming a cascaded pipeline.
+Leverage the following hierarchical probabilistic decomposition via marginalization:
 $p(\boldsymbol{u}|\boldsymbol{y})=\int p(\boldsymbol{u},\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}=\int p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}=\mathbb{E}_{\boldsymbol{m}\sim p(\boldsymbol{m}|\boldsymbol{y})}[\,p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,],$ (1)
-where $\boldsymbol{m}$ represents the principal structure (the large-scale component) of $\boldsymbol{u}$.
-This formulation naturally decomposes the task into two subtasks: modeling $p(\boldsymbol{m}|\boldsymbol{y})$ to infer the principal structure from sparse observations, and modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ to reconstrut the full field $\boldsymbol{u}$ by adding high-frequency details to $\boldsymbol{m}$ based on both $\boldsymbol{m}$ and $\boldsymbol{y}$.
-Since the principal structures of the signal are often sparse, reconstructing $\boldsymbol{m}$ from sparse observations—i.e., modeling $p(\boldsymbol{m}|\boldsymbol{y})$—is generally well-posed.
-Furthermore, compared to directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$, modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ incorporates an additional condition, constraining the reconstruction of $\boldsymbol{u}$ to remain consistent with the principal structures $\boldsymbol{m}$.
+where $\boldsymbol{m}$ represents the coarse-scale structure (the coarse-scale component) of $\boldsymbol{u}$.
+This formulation naturally decomposes the task into two subtasks: modeling $p(\boldsymbol{m}|\boldsymbol{y})$ to infer the coarse-scale structure $\boldsymbol{m}$ from sparse observations $\boldsymbol{y}$; and modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ to further reconstrut the full field $\boldsymbol{u}$ by adding refined details to $\boldsymbol{m}$, conditioned on both $\boldsymbol{m}$ and $\boldsymbol{y}$.
+At coarse scales, physical fields are dominated by low-frequency, smoothly varying components governed by global constraints, which can be represented with far fewer degrees of freedom than fine-scale fluctuations.
+As a result, reconstructing $\boldsymbol{m}$ from sparse observations –i.e., modeling $p(\boldsymbol{m}|\boldsymbol{y})$ – is generally well-posed.
+Furthermore, compared to directly modeling $p(\boldsymbol{u}|\boldsymbol{y})$, modeling $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ incorporates an additional condition, constraining the reconstruction of $\boldsymbol{u}$ to remain consistent with the coarse-scale structure $\boldsymbol{m}$.
 Consequently, the degree of ill-posedness is substantially alleviated through this cascaded pipeline.
@@
 Then, a conditional diffusion model trained with the proposed mask-cascade training strategy to approximate $\mathbb{E}_{\boldsymbol{m}\sim p(\boldsymbol{m}|\boldsymbol{y})}[\,p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,]$.
-Specifically, once the autoencoder is trained and fixed, random sparse measurements $\boldsymbol{y}$ are generated from $\boldsymbol{u}$ and mapped through the autoencoder to produce corresponding $\boldsymbol{m}$.
+Specifically, once the autoencoder is trained and fixed, random sparse measurements $\boldsymbol{y}$ are mapped through the autoencoder to produce corresponding $\boldsymbol{m}$.
 The conditional diffusion model is then trained to generate $\boldsymbol{u}$ conditioned on both $\boldsymbol{m}$ and $\boldsymbol{y}$, effectively learning to reconstruct the full field from varying sparse observations.
-2.2 Functional autoencoder for principal structure reconstruction
-In practical engineering, sensors are often randomly placed, and many physical fields involve complex geometric boundaries, making field reconstruction from unstructured data challenging [chai2020deep].
+2.2 Functional autoencoder for coarse-scale structure reconstruction
+Figure 2: The masked training scheme of functional autoencoder.
+The trained functional autoencoder is to approximate $p(\boldsymbol{m}|\boldsymbol{y})$ with $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$ as it learns the deterministic mapping between $\boldsymbol{y}$ and $\boldsymbol{m}$.
+$x_{e}$ and $u_{e}$ denote the coordinates and field values fed into the encoder, while $x_{d}$ and $u_{d}$ represent the coordinates and reconstructed values produced by the decoder.
+In practical engineering, sensors are often randomly placed (or some sensors are even movale), and many physical fields involve complex geometric boundaries, making field reconstruction from unstructured data challenging [chai2020deep].
 To address this, we adopt a neural operator-based functional autoencoder [bunker2024autoencoders], which enables coordinate-based input for neural networks [regazzoni2024learning, krishnapriyan2023learning].
-The functional autoencoder extracts latent features from unstructured data and robustly reconstruct the principal structure from random and sparse inputs.
-We model the inverse mapping from sparse observations $\boldsymbol{y}$ to the principal structure $\boldsymbol{m}$ using a deterministic autoencoder rather than a probabilistic model (e.g., a variational autoencoder).
-This choice is justified by the low uncertainty inherent in reconstructing $\boldsymbol{m}$ from $\boldsymbol{y}$ due to: (1) the strong sparse prior of $\boldsymbol{m}$ constraining the solution space since large-scale patterns vary smoothly and can be represented with few basis functions; (2) average pooling operation in the autoencoder promoting invariant and robust feature learning; and (3) our masked training strategy explicitly enhancing model’s robustness to noise and varying sensing patterns in $\boldsymbol{y}$.
-As a result, the posterior $p(\boldsymbol{m}|\boldsymbol{y})$ is highly concentrated, allowing us to approximate it as $p(\boldsymbol{m}|\boldsymbol{y})\approx\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$, where $\delta(\cdot)$ is a Dirac delta function and $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is the autoencoder output.
+The functional autoencoder extracts latent features from unstructured data and robustly reconstruct the coarse-scale structure from random and sparse inputs.
+We model the inverse mapping from sparse observations $\boldsymbol{y}$ to the coarse-scale structure $\boldsymbol{m}$ using a deterministic autoencoder rather than a probabilistic model (e.g., a variational autoencoder).
+This choice is justified by the low uncertainty inherent in reconstructing $\boldsymbol{m}$ from $\boldsymbol{y}$ due to: (1) the strong sparse prior of $\boldsymbol{m}$ constraining the solution space since coarse-scale patterns vary smoothly and can be represented with few basis functions; (2) an average pooling operation in the autoencoder promoting invariant and robust feature learning; and (3) our masked training strategy explicitly enhancing model’s robustness to noise and varying sensing patterns in $\boldsymbol{y}$.
+As a result, the posterior $p(\boldsymbol{m}|\boldsymbol{y})$ is highly deterministic, allowing us to approximate it as:
+$p(\boldsymbol{m}|\boldsymbol{y})\approx\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$ (2)
+where $\delta(\cdot)$ is a Dirac delta function and $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is the autoencoder output.
 This avoids instability from learning a full posterior variance, and still achieves accurate and robust reconstruction, as confirmed in our experiments.
@@
 $\{(x_{i},u(x_{i}))|i=1,2,...I\}\rightarrow\rho(\text{pool}(\{\kappa(x_{i},u(x_{i});\zeta_{1})|i=1,2,...I\};\zeta_{2}))$ (4)
-where pool is a pooling operation invariant to the order of its inputs—in our case, the average pooling.
-The average pooling operation reduces the discrepancy between latent representations derived from sparse and dense inputs, thereby endowing the functional autoencoder with the capability for accurate sparse-data reconstruction.
+where pool denotes a pooling operation that is invariant to the order of its inputs—specifically, average pooling in our implementation.
+Owing to its permutation-invariant property, average pooling enables the functional autoencoder to process arbitrarily ordered and distributed sparse inputs.
+Moreover, it can adaptively attenuating high-frequency variations while preserving low-frequency components, which is consistent with the well-known smoothing effect of averaging operations.
+Since average pooling reduces the discrepancy between latent representations obtained from sparse and dense observations, the learned latent space becomes largely insensitive to the input sampling ratio, endowing the functional autoencoder with robust and accurate reconstruction capability under extreme sparsity.
 The decoder $\mathcal{D}$ maps a latent vector $z\in\mathcal{Z}$ back to a function $\mathcal{D}(z;\psi):\Omega\rightarrow\mathbb{R}^{m}$.
@@
 […]
+where the second term, $\mathcal{E}(u;\zeta)||_{2}^{2}$ with coefficient $\beta$, plays the role of regularisation, ensuring that the latent representation is smooth and structured.
 […]
-[…]
-2.3 Conditional diffusion model for detail generation
-After reconstructing the principal structures with the functional autoencoder, a conditional diffusion model is trained to generate the small-scale components, defined as the difference between the ground truth and the reconstructed principal structures, conditioned on the principal structures.
-By defining the detail or residual $\boldsymbol{d}=\boldsymbol{u}-\boldsymbol{m}$, learning $p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})$ is equivalent to learning $p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})$ as it is a deterministic relationship.
-Modeling the residual $\boldsymbol{d}$ simplifies the learning task, reduces complexity of modeling, and improves training stability and convergence.
-It also introduces an inductive bias, constraining the reconstructed $\boldsymbol{u}$ to remain close to $\boldsymbol{m}$, thereby restricting the solution space and mitigating ill-posedness.
-As discussed above, the functional autoencoder approximates $p(\boldsymbol{m}|\boldsymbol{y})$ with a Dirac delta function $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
-Consequently, the conditional diffusion model should be trained to model:
-$p(\boldsymbol{u}|\boldsymbol{y})=\int p(\boldsymbol{u}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}$ $=\int p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}|\boldsymbol{y})\,d\boldsymbol{m}\approx\int p(\boldsymbol{d}|\boldsymbol{m},\boldsymbol{y})\,\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))\,d\boldsymbol{m}=p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y}).$
+2.3 Conditional diffusion models for refined-scale detail generation
+In Cas-Sensing, the second stage is designed to generate refined-scale details conditioned on the recovered coarse-scale structures, while explicitly enforcing consistency with sparse measurements.
+To this end, we adopt a Denoising Diffusion Probabilistic Model (DDPM) as the generative backbone, owing to its Bayesian interpretation of the sampling process, which enables additional conditioning constraints to be injected at inference time without retraining the model.
+Specifically, the reverse diffusion process can be viewed as sampling from a data-driven prior, and measurement consistency can be incorporated by reformulating unconditional sampling as posterior sampling under an observation likelihood.
+This property allows sparse observational constraints to be enforced through gradient-based corrections during denoising (e.g., manifold-constrained gradients), achieving data fidelity while preserving the learned data manifold, and making DDPMs particularly suitable for deployment under varying sensor configurations and sparsity levels.
+After reconstructing the coarse-scale structures $\boldsymbol{m}$ with the functional autoencoder, the conditional diffusion model is trained to generate the refined-scale components defined as the residual $\boldsymbol{d}=\boldsymbol{u}-\boldsymbol{m}$.
+Under this deterministic decomposition, learning the conditional distribution $p(\boldsymbol{u}\mid\boldsymbol{m},\boldsymbol{y})$ is equivalent to learning $p(\boldsymbol{d}\mid\boldsymbol{m},\boldsymbol{y})$.
+Modeling the residual significantly simplifies the learning task by reducing distributional complexity, improving training stability and convergence, and introducing an inductive bias that constrains the reconstructed field to remain close to the coarse-scale estimate.
+As discussed in Eq.
+(2), the functional autoencoder approximates $p(\boldsymbol{m}\mid\boldsymbol{y})$ with a Dirac delta distribution $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$, leading to the following cascaded probabilistic formulation:
+$p(\boldsymbol{u}\mid\boldsymbol{y})=\int p(\boldsymbol{u}\mid\boldsymbol{m},\boldsymbol{y})\,p(\boldsymbol{m}\mid\boldsymbol{y})\,d\boldsymbol{m}\approx\int p(\boldsymbol{d}\mid\boldsymbol{m},\boldsymbol{y})\,\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))\,d\boldsymbol{m}=p(\boldsymbol{d}\mid\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y}).$ (7)
 2.3.1 Conditional denoising diffusion probabilistic model
-Training of denoising diffusion probabilistic models (DDPMs) consists of a diffusion process and a denoising process.
+Training of DDPMs consists of a diffusion process and a denoising process.
 Diffusion process transforms $\boldsymbol{d}_{0}$ from the real data distribution $\boldsymbol{d}_{0}\sim p(\boldsymbol{d}_{0})$ into a pure Gaussian noise $\boldsymbol{d}_{t}\sim\mathcal{N}(0,\boldsymbol{I})$ by successively applying the following Markov diffusion kernel:
@@
 For sampling, $\boldsymbol{d}_{t}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{d}_{0}+(1-\bar{\alpha}_{t})\boldsymbol{\epsilon},\ \boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
-Given $\boldsymbol{d}_{t}$, the denoising process aims to recover $\boldsymbol{d}_{0}$ by recursively learning the transition from $\boldsymbol{d}_{t-1}$ to $\boldsymbol{d}_{t}$, which is defined as the following Gaussian distribution:
+Reversely, given $\boldsymbol{d}_{t}$, the denoising process aims to recover $\boldsymbol{d}_{0}$ by recursively learning the transition from $\boldsymbol{d}_{t-1}$ to $\boldsymbol{d}_{t}$, which is defined as the following Gaussian distribution:
 $q_{\theta}(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t})=\mathcal{N}\left(\boldsymbol{d}_{t-1};\boldsymbol{\mu}_{\theta}(\boldsymbol{d}_{t},t),\boldsymbol{\Sigma}_{\theta}(\boldsymbol{d}_{t},t)\right),$ (10)
@@
 However, this unconditional scheme fails to preserve the uniqueness of reconstruction due to the lack of guidance during generation.
-Therefore, $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is injected as the condition for the denoising process:
+Therefore, the coarse-scale inference $\hat{\boldsymbol{m}}(\boldsymbol{y})$ is injected as the condition for the denoising process:
 $q_{\theta}(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))=\mathcal{N}(\boldsymbol{d}_{t-1};\boldsymbol{\mu}_{\theta}(\boldsymbol{d}_{t},t,\hat{\boldsymbol{m}}(\boldsymbol{y})),\sigma^{2}_{t}\mathbf{I}),$ (11)
@@
 This process is repeated until $t=1$, at which point $\boldsymbol{d}_{0}$ is taken as the final reconstructed sample.
-2.3.2 Manifold constrained gradient
+2.3.2 Training stage: Mask-cascade training
+Figure 3: The proposed mask-cascade training (MCT) of the conditional diffusion model.
+A key challenge is that the functional autoencoder, when provided with different sparse samplings from the same field, reconstructs slightly different coarse-scale structures $\hat{\boldsymbol{m}}(\boldsymbol{y})$.
+Since it is infeasible to enumerate all possible reconstructions for training, directly using a fixed dataset would restrict the generalization ability of the diffusion model.
+In this regard, we propose a mask-cascade training strategy to improve the robustness of the conditional diffusion model under varying sparsity patterns.
+Our strategy explicitly integrates the stochasticity of sparse sampling into training the conditional diffusion model.
+As illustrated in Fig.
+3, once the functional autoencoder is fully trained and its parameters are frozen, the training of the conditional diffusion model proceeds as follows: in each training step, we randomly generate a sparse mask with fixed mask ratio and apply it to the ground-truth field $\boldsymbol{u}$ to obtain a partial observation $\boldsymbol{y}$.
+Notably, the mask ratio here is chosen to be significantly lower than that used for masked training the functional autoencoder, so as to explicitly train the conditional diffusion model to operate under extremely sparse observations.
+This observation is then passed through the trained (fixed) functional autoencoder to reconstruct an approximate coarse-scale structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$.
+The conditional diffusion model is then conditioned on $\hat{\boldsymbol{m}}(\boldsymbol{y})$, and trained to generate the fine-scale component $\boldsymbol{d}=\boldsymbol{u}-\hat{\boldsymbol{m}}(\boldsymbol{y})$.
+By repeatedly implementing different masks across training steps, the conditional diffusion model is continuously exposed to a diverse ensemble of approximate coarse-scale structure reconstructions.
+This effectively augments the training distribution of conditioning inputs without the need for explicitly precomputing or storing all possible variants.
+As a result, the conditional diffusion model learns to model the conditional distribution $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$, thereby achieving improved robustness and consistency across varying sparsity patterns.
+2.3.3 Inference stage: Manifold constrained gradient
 The conditional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by training the conditional diffusion model.
-The $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ should be modeled to further preserve the uniqueness of reconstruction and impose measurement consistency.
+According to the formulation in Eq.(7), the $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ should be modeled to further preserve the uniqueness of reconstruction and impose measurement consistency.
 However, directly model the condtional distribution $p(\boldsymbol{d}_{t-1}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ is impossible due to the immense variability of sparse measurement patterns.
-We model it during inference instead of training of conditional diffusion model.
+We model it during the inference stage, instead of training of such a conditional diffusion model.
 Specifically, the Manifold Constrained Gradient (MCG) [chung2022improving] is used to inject an additional gradient term to guide the generation towards observed measurements without retraining the model.
 This gradient refers to the score function in score-based generative models (SGMs) [song2020score], defined as the gradient of the log-density of the data distribution $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
-Although DDPMs and SGMs were originally proposed under different formulations—predicting added noise in DDPMs and estimating score functions in SGMs—it has been shown that they are theoretically equivalent under certain parameterizations [song2020score].
+Although DDPMs and SGMs were originally proposed under different formulations – predicting added noise in DDPMs and estimating score functions in SGMs – it has been shown that they are theoretically equivalent under certain parameterizations [song2020score].
 Specifically, when a DDPM is trained to predict the Gaussian noise $\boldsymbol{\epsilon}_{\theta}$ added at each step, the network implicitly learns the score function of the perturbed data distribution $p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ up to a scaling factor:
 $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))\approx-\frac{1}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\boldsymbol{d}_{t},t,\hat{\boldsymbol{m}}(\boldsymbol{y})),$ (13)
-Therefore, the sampling equation in Eq. 10 can be reformulated as:
-$\mathbf{\boldsymbol{d}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{\boldsymbol{d}}_{t}+(1-\alpha_{t})\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))\right)+\sigma_{t}\boldsymbol{\epsilon},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (12)
 This connection enables MCG to be directly applied to DDPMs.
@@
 where $\nabla_{\boldsymbol{d}_{t}}\log p_{t}(\boldsymbol{d}_{t}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$ has been modeled by conditional diffusion model.
-To approximate the likelihood term $p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$, we assume that it follows a Gaussian form, i.e.,
-$p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))\propto\exp\left(-\frac{1}{2\sigma_{c}^{2}}\|\boldsymbol{y}-\mathcal{M}(\hat{\boldsymbol{m}}(\boldsymbol{y})+\boldsymbol{d}_{t})\|_{2}^{2}\right),$ (14)
-where $\mathcal{M}(\cdot)$ denotes the mask operator and $\sigma_{c}^{2}$ denotes the variance associated with measurement noise.
-$\sigma_{c}^{2}$ is set as 10000 in all experiments of this study.
+To approximate the likelihood term $p_{t}(\boldsymbol{y}|\boldsymbol{d}_{t},\hat{\boldsymbol{m}}(\boldsymbol{y}))$, we assume that it follows a Gaussian form.
 Under this assumption, its gradient with respect to $\boldsymbol{d}_{t}$ yields:
 […]
-Therefore, the sampling formula without retraining the model is given as:
-[…]
+where $\mathcal{M}(\cdot)$ denotes the mask operator and $\sigma_{c}^{2}$ denotes the variance associated with measurement noise. $\sigma_{c}^{2}$ is set as 10000 in all experiments of this study.
+According to the Bayes rule in Eq.
+14, the measurement consistency can be injected into the trained diffusion model by adding this new gradient into the sampling formulation in Eq.
+12.
 However, it drives the intermediate states at each denoising step to match the observations, which pulls the sample path away from the learned data manifold.
@@
 […]
-2.3.3 Mask-cascade training
-We propose a mask-cascade training strategy to improve the robustness of the conditional diffusion model under varying sparsity patterns.
-A key challenge is that the functional autoencoder, when provided with different sparse samplings of the same field, reconstructs slightly different principal structures.
-Since it is infeasible to enumerate all possible reconstructions for training, directly using a fixed dataset would restrict the generalization ability of the diffusion model.
-To address this issue, our strategy explicitly integrates the stochasticity of sparse sampling into training the conditional diffusion model.
-As illustrated in Fig.
-1(b), once the functional autoencoder is fully trained and its parameters are frozen, the training of the conditional diffusion model proceeds as follows: in each training step, we randomly generate a sparse mask and apply it to the ground-truth field $\boldsymbol{u}$ to obtain a partial observation $\boldsymbol{y}$.
-This observation is then passed through the functional autoencoder to reconstruct an approximate principal structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$.
-The conditional diffusion model is then conditioned on $\hat{\boldsymbol{m}}(\boldsymbol{y})$, and trained to generate the fine-scale component $\boldsymbol{d}=\boldsymbol{u}-\hat{\boldsymbol{m}}(\boldsymbol{y})$.
-By repeatedly sampling different masks across training steps, the conditional diffusion model is continuously exposed to a diverse ensemble of approximate principal structure reconstructions.
-This effectively augments the training distribution of conditioning inputs without the need for explicitly precomputing or storing all possible variants.
-As a result, the conditional diffusion model learns to model the conditional distribution $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$, thereby achieving improved robustness and consistency across varying sparsity patterns.
-As shown in Fig.1(c), the inference process of autoencoder-diffusion cascade proceeds as follows.
-Given arbitrary sparse observations $\boldsymbol{y}$, the functional autoencoder first reconstructs the principal structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$.
+As shown in Fig.4, the inference process of autoencoder-diffusion cascade proceeds as follows.
+Given arbitrary sparse observations $\boldsymbol{y}$, the functional autoencoder first reconstructs the coarse-scale structure $\hat{\boldsymbol{m}}(\boldsymbol{y})$.
 This reconstructed structure is then provided as a condition to the conditional diffusion model, which generates the fine-scale details $\hat{\boldsymbol{d}}$ through the manifold constrained gradient-enhanced sampling scheme in Eq.
@@
 It is worth noting that, during inference, Cas-Sensing can still achieve reliable reconstruction even when the sampling rate of sparse observations differs significantly from the rates used in mask-cascade training.
-(a) The masked training scheme of functional autoencoder.
-The trained functional autoencoder is to approximate $p(\boldsymbol{m}|\boldsymbol{y})$ with $\delta(\boldsymbol{m}-\hat{\boldsymbol{m}}(\boldsymbol{y}))$ as it learns the deterministic mapping between $\boldsymbol{y}$ and $\boldsymbol{m}$.
-$x_{e}$ and $u_{e}$ denote the coordinates and field values fed into the encoder, while $x_{d}$ and $u_{d}$ represent the coordinates and reconstructed values produced by the decoder.
-(b) The mask-cascade training scheme of conditional diffusion model, which ensure the conditional diffusion model’s accurate capturing $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}))$.
-(c) The inference scheme of Cas-Sensing based on an autoencoder-diffsusion cascade, which is the process of sampling from $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$.
-Manifold Constrained Gradient (MCG) is used to impose measurement consistency and sample from $p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$ without retraining the model.
-Figure 1: Overview of the proposed Cascaded Sensing (Cas-Sensing) based on an autoencoder-diffusion cascade.
+Figure 4: The inference scheme of Cascaded Sensing.
+Given different sparse measurements of the same target physical field, the framework samples from the conditional distribution $p(\boldsymbol{u}|\boldsymbol{y})\approx p(\boldsymbol{d}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$, producing multiple reconstruction samples.
+Measurement consistency is imposed with Manifold Constrained Gradient (MCG) during diffusion sampling, without retraining the model.
 3 Results
 3.1 Reconstructing cylinder flow velocity fields with simulation data
-To evaluate the effectiveness of the proposed Cas-Sensing based on an autoencoder-diffusion cascade in reconstructing full-field data from randomly sparse observations under varying geometric boundary conditions, we first constructed a 2D cylinder flow dataset.
+To evaluate the effectiveness of the proposed Cas-Sensing in reconstructing full-field data from randomly sparse observations under varying geometric boundary conditions, we first constructed a synthetic 2D cylinder flow dataset.
 This dataset captures incompressible fluid dynamics around a 2D circular cylinder within a channel:
@@
 $\frac{\partial\mathbf{u}}{\partial t}+(\mathbf{u}\cdot\nabla)\mathbf{u}$ $=\mu\nabla^{2}\mathbf{u}-\frac{1}{\rho}\nabla p.$ (20)
-Specifically, the computational domain is defined as a rectangular region of size $1.6\times 0.6$, within which both the cylinder center and radius are randomly sampled.
+Specifically, the computational domain is defined as a rectangular region of size $1.6\times 0.6\ m$, within which both the cylinder center position and radius value are randomly sampled.
 This procedure yields a dataset comprising 96 distinct boundary configurations, each associated with 100 temporal snapshots.
@@
 The visualization of a sample from the Cylinder Flow is presented in Fig.
-2.
+5.
 For the purpose of model training and evaluation, the dataset is partitioned based on boundary configurations rather than temporal snapshots, ensuring that the test set contains unseen geometrical variations.
@@
 Specifically, the 32-dimensional latent vectors are projected into a two-dimensional space using the t-distributed Stochastic Neighbor Embedding (t-SNE) technique [maaten2008visualizing].
-Each point in Fig.
+Each scattered point in Fig.
 6 corresponds to the latent variable of a single flow snapshot.
-It can be observed that the training data associated with different internal boundary configurations are well separated in the latent space, and the latent vectors of snapshots from the same configuration form circular patterns.
+It can be observed that the training data associated with different internal boundary configurations are well separated in the latent space, while the latent vectors of snapshots from the same configuration form circular patterns.
 This circular structure reflects the inherent periodicity of the velocity fields in cylinder wake flows.
-Moreover, the latent vectors corresponding to the test dataset exhibit the same characteristic structures, indicating that the functional autoencoder has effectively captured both the temporal and geometric features of the flow fields.
+Moreover, the latent vectors corresponding to the test dataset exhibit the same characteristic structures, indicating that the functional autoencoder has effectively captured both the temporal and geometric features of the given flow fields.
 Figure 6: t-SNE visualization of latent representations learned by the functional autoencoder for cylinder flow fileds.
@@
 The distributions of the resulting reconstruction errors, quantified by the root mean square error (RMSE), are illustrated in Fig.
-4.
+7.
 Quantitative results confirm the strong robustness of the functional autoencoder in reconstructing cylinder wake flows under varying levels of input sparsity.
@@
 These results demonstrate that the autoencoder is capable of faithfully recovering the dominant flow features even with extremely sparse measurements, underscoring its resilience to data sparsity and strong generalization ability.
-Figure 4: The reconstructed principal structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
+Figure 7: The reconstructed coarse-scale structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
 After training the functional autoencoder, we freeze its parameters and employ it to provide structural conditions for a conditional diffusion model.
@@
 Representative results, shown in Fig.
-5(a), indicate that the generated fields reproduce the ground truth with high accuracy, capturing both the periodic vortex shedding and the influence of varying boundary configurations.
+8(a), indicate that the generated fields reproduce the ground truth with high accuracy, capturing both the periodic vortex shedding and the influence of varying boundary configurations.
 To further probe reconstruction reliability, we select a test snapshot and impose random masks with different sampling ratios ($0.1\%,0.5\%,1\%,3\%$, and $5\%$).
 For each ratio, 100 independent reconstructions are carried out, and the kernel density distributions of RMSE are computed, as presented in Fig.
-5(b).
+8(b).
 The results demonstrate a clear trend: at extremely sparse ratios such as $0.1\%$, the reconstructions remain accurate but exhibit slightly higher variance; as the ratio increases to $1\%$, both the mean error and uncertainty decrease significantly, reflecting enhanced stability of the generated fields.
@@
 The example of stereo-image pair process by Wave Acquisition Stereo System (WASS) is presented in Fig.
-6.
+9.
 The data set used in this study is Acqua Alta stereo data set including 8000 image frames, recording a half hour sea evaluation with an imaging system installed on the north-east side of the Acqua Alta oceanographic research tower.
 Figure 9: An example of stereo-image pair (from the Acqua Alta tower in the northern Adriatic Sea, Italy) and 3-D wave field on top of the right image (the colorscale is proportional to the sea surface height)[guimaraes2020data].
-First, we train a functional autoencoder to reconstruct the principal structures of sea surface wave fields using 7000 image frames, with the remaining 1000 frames held out as a test set.
+First, we train a functional autoencoder to reconstruct the coarse-scale structures of sea surface wave fields using 7000 image frames, with the remaining 1000 frames held out as a test set.
 During masked training, $50\%$ of the data points in each image are provided to the encoder as input for feature extraction, while the model is trained to reconstruct the complementary $50\%$ of the points.
@@
 Each masked input is then encoded and decoded on the full grid, and kernel density estimates of the reconstruction root mean square error (RMSE) are computed, as presented in Fig.
-7.
+10.
 From Fig.
-7, it is evident that when the input ratio is above approximately $3\%$, the functional autoencoder achieves highly accurate and stable reconstructions of the principal structures, with RMSE means remaining close to 0.40 and variances below 0.015.
+10, it is evident that when the input ratio is above approximately $3\%$, the functional autoencoder achieves highly accurate and stable reconstructions of the coarse-scale structures, with RMSE means remaining close to 0.40 and variances below 0.015.
 As the input ratio decreases further, the reconstruction quality gradually degrades.
 At $1\%$ input, for example, the mean RMSE increases to about 0.43 and the variance nearly doubles, and at $0.5\%$ input the mean RMSE reaches approximately 0.45 with the largest spread in the error distribution.
-Despite this deterioration, the reconstructions still preserve the correct large-scale patterns of the wave field, highlighting the strong robustness of the functional autoencoder to severe data sparsity in capturing the dominant structures.
-Figure 7: The reconstructed principal structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
-When the functional autoencoder is trained and fixed, it is used to output principal structures according to sparse measurements as conditions to train a conditional diffusion model with the proposed mask-cascade training strategy.
+Despite this deterioration, the reconstructions still preserve the correct coarse-scale patterns of the wave field, highlighting the strong robustness of the functional autoencoder to severe data sparsity in capturing the dominant structures.
+Figure 10: The reconstructed coarse-scale structures and kernel density estimates for RMSE on the reference sample across 100 randomly chosen meshes.
+When the functional autoencoder is trained and fixed, it is used to output coarse-scale structures according to sparse measurements as conditions to train a conditional diffusion model with the proposed mask-cascade training strategy.
 The input point ratio of the fixed functional autoencoder is set as $0.5\%$ to enhance the robustness of conditional diffusion model to sparsity.
-To test the generation results, given $0.5\%$ randomly selected observation points from each test sample, the functional autoencoder reconstructs the corresponding principal structures.
-These structures are then used as conditions to input the conditional diffusion model while manifold constrained gradients are incorporated into the sampling process to ensure consistency between the generated full-field data and the observed $0.5\%$ points.
+To test the generation results, given $0.5\%$ randomly selected observation points from each test sample, the functional autoencoder reconstructs the corresponding coarse-scale structures.
+These structures are then used as conditions to input the conditional diffusion model, while manifold constrained gradients are incorporated into the sampling process to ensure consistency between the generated full-field data and the observed $0.5\%$ points.
 As shown in Fig.
-8(a), the generated fields exhibit high fidelity to the ground truth.
+11(a), the generated fields exhibit high fidelity to the ground truth.
 It is important to note that each generated realization corresponds to a single sample from $p(\boldsymbol{u}|\hat{\boldsymbol{m}}(\boldsymbol{y}),\boldsymbol{y})$.
@@
 The results, shown in Fig.
-8, where the mean field of generations match the ground truth very well.
+11, where the mean field of generations match the ground truth very well.
 As the percentage of observed points increases from $0.2\%$ to $1\%$, the mean RMSE decreases notably, indicating improved reconstruction quality with denser observations.
 In particular, the drop in both mean error and variance around the $1\%$ input level suggests that even a relatively small increase in observations can significantly stabilize the generation.
-When the ratio is further increased to $3\%$ and $5\%$, the mean RMSE continues to decrease slightly, approaching a lower bound, while the variance remains at a low level since the principal structures reconstruction results are more stable under these ratios, reflecting both accuracy and robustness.
+When the ratio is further increased to $3\%$ and $5\%$, the mean RMSE continues to decrease slightly, approaching a lower bound, while the variance remains at a low level since the coarse-scale structures reconstruction results are more stable under these ratios, reflecting both accuracy and robustness.
 It is worth noting that although the conditional diffusion model is trained with $0.5\%$ input points, it generalizes well to different input ratios and consistently delivers satisfactory reconstructions even with only $0.2\%$ points.
@@
 The downsampled fields were then evenly partitioned into 15 subregions, each with a resolution of 480 × 128, as illustrated in Fig.
-9, forming SST dataset.
+12, forming SST dataset.
 For model training and evaluation, the first $80\%$ of the data were used as the training set, and the remaining $20\%$ were reserved for testing.
@@
 As illustrated in Fig.
-10(a) (Condition), the trained functional autoencoder is capable of accurately reconstructing the large-scale background patterns from only partial observations, capturing the essential spatial variability of the ocean surface temperature fields.
+13(a) (Condition), the trained functional autoencoder is capable of accurately reconstructing the coarse-scale background patterns from only partial observations, capturing the essential spatial variability of the ocean surface temperature fields.
 Once the autoencoder is fully trained, it is frozen and integrated into the Cascaded Sensing framework to provide real-time background field reconstructions that serve as conditioning information for the conditional diffusion model during mask-cascade training.
@@
 A representative reconstruction (RMSE = 0.0144) is shown in Fig.
-10(a).
-Remarkably, as evident in the zoomed-in views, the Cascaded Sensing framework not only reconstructs the global background temperature field with high accuracy but also successfully captures fine-scale turbulent structures, which are typically difficult to recover from such sparse observations.
+13(a).
+Remarkably, as evident in the zoomed-in views, the Cascaded Sensing framework not only reconstructs the global background temperature field with high accuracy but also successfully captures fine-scale turbulent structures, which are typically highly challenging to recover from such sparse observations.
 To systematically assess the robustness of the proposed approach, we further conduct experiments under a range of observation ratios ($5\%,3\%,1\%,0.5\%,0.1\%$).
@@
 The resulting mean and standard deviation of RMSE are summarized in Fig.
-10(b).
+13(b).
 As expected, both the mean and standard deviation of RMSE increase as the observation ratio decreases.
@@
 Figure 13: The full-field reconstruction results of global sea temperature fields.
-4 Discussion
-In this study, we conduct a comprehensive evaluation of the proposed Cascaded Sensing (Cas-Sensing) framework under a broad range of scenarios characterized by different sensor configurations, sampling sparsities, and geometric boundaries.
-Our results clearly show that the hierarchical structure of Cas-Sensing can substantially mitigate the ill-posedness of reconstructing multi-scale physical fields from extremely sparse measurements.
-By decomposing the reconstruction process into two stages—large-scale structure recovery via a functional autoencoder and fine-scale detail generation through a conditional diffusion model—the proposed method achieves accurate and stable reconstructions with high data fidelity.
-A central strength of this cascaded design lies in its ability to reduce the reconstruction ambiguity caused by sparse and irregular inputs.
-By first establishing reliable coarse-scale conditioning, the diffusion model no longer needs to infer the entire field from incomplete observations, but focuses on capturing the missing small-scale structures.
-This hierarchical formulation enhances reconstruction stability and reduces the variability across different sampling patterns.
-Moreover, the mask-cascade training strategy significantly improves the model’s robustness and generalization, enabling it to adapt to diverse sensor layouts without the need for retraining.
-These advantages make Cas-Sensing particularly effective for scientific applications.
-For example, in fluid dynamics, accurate flow field reconstruction from sparse sensors is crucial for monitoring and control; in climate and ocean science, reconstructing global or regional temperature fields from sparse measurements enables more reliable environmental assessment; and in geophysical applications, the ability to generalize across sampling configurations supports practical deployment in complex field settings.
-Overall, Cas-Sensing provides a flexible and efficient framework for reconstructing multi-scale fields.
-By leveraging hierarchical representation learning and generative modeling, it delivers both accuracy and robustness across a variety of reconstruction scenarios, offering a promising pathway for advancing data-driven scientific sensing and prediction.
-5 Conclusion and future work
-In this study, we have presented Cascade Sensing (Cas-Sensing) as an effective framework for reconstructing full-field data from highly sparse observations.
-By leveraging an autoencoder-diffusion cascade framework, Cas-Sensing is capable of producing accurate and coherent reconstructions even when the input measurements are extremely limited.
-The mask-cascade training strategy, combined with manifold constrained gradient, enhances the robustness of the model to diverse sparse conditions and ensures the consistency of the reconstructed physical fields.
-These results demonstrate the potential of Cas-Sensing as a powerful tool for efficient, data-driven sensing and predictive modeling in complex physical systems.
-Nevertheless, several limitations remain in the current implementation.
-Firstly, the present study focuses exclusively on two-dimensional spatial domains, leaving the extension to higher-dimensional settings unexplored.
-Scaling Cas-Sensing to three-dimensional or higher-dimensional problems may introduce additional computational and modeling challenges that warrant careful investigation.
-Secondly, the current framework addresses only spatial reconstruction, without explicitly modeling temporal dynamics.
-Consequently, the method cannot fully capture spatiotemporal evolution in dynamic systems, which may limit its applicability in real-time monitoring or predictive scenarios.
-Thirdly, the generative component currently relies on a vanilla DDPM, which, while effective, may impose constraints on sampling speed and computational efficiency.
-Alternative generative models, such as flow matching, could potentially accelerate sampling and further improve reconstruction fidelity.
-Future research should aim to overcome these limitations.
-Extending Cas-Sensing to higher-dimensional domains would broaden its applicability to more complex physical systems, including 3D flow fields or volumetric sensor networks.
-Integrating temporal modeling within the cascade framework could enable spatiotemporal reconstruction, capturing dynamic evolution in addition to spatial correlations, and allowing the framework to support real-time prediction tasks.
-Moreover, the adoption of more advanced generative models could enhance sampling efficiency without sacrificing accuracy, making Cas-Sensing more practical for large-scale deployments.
-Collectively, these developments would not only enhance the versatility and scalability of Cas-Sensing but also reinforce its potential as a robust and reliable tool for data-driven sensing, reconstruction, and prediction in complex physical environments.
+4 Discussion and conclusion
+In this study, we presented and systematically evaluated Cascaded Sensing (Cas-Sensing) as a probabilistic paradigm for reconstructing multi-scale physical fields under extreme data sparsity.
+By explicitly recognizing sparse sensing as a severely ill-posed inverse problem with intrinsic non-uniqueness, Cas-Sensing reformulates full-field reconstruction as a hierarchical probabilistic inference task, rather than attempting to learn a direct deterministic mapping from sparse observations to full fields.
+Extensive experiments across a wide range of sensor configurations, sampling sparsities, and geometric boundaries consistently demonstrate that this hierarchical formulation effectively mitigates the ill-posedness inherent in sparse reconstruction problems.
+A central strength of the proposed framework lies in the introduction of an explicit intermediate representation that captures coarse-scale structures.
+By first recovering physically meaningful coarse-scale components and geometric features via a functional autoencoder, the original ill-posed problem is decomposed into two substantially better-conditioned subproblems.
+This decomposition constrains subsequent reconstruction to remain consistent with dominant structures, allowing the conditional diffusion model to focus exclusively on generating missing refined-scale details as residuals.
+As a result, reconstruction responsibilities are decoupled across spatial scales, improving stability and fidelity while avoiding information loss associated with unified latent-space generation.
+The robustness and generalization capabilities of Cas-Sensing are further enhanced by the proposed mask-cascade training strategy and manifold-constrained gradient-based sampling.
+By exposing the diffusion model to a distribution of imperfect conditioning structures induced by varying sparsity patterns, the model learns a conditional distribution that remains stable across diverse sensor layouts and extreme sparsity levels.
+During inference, measurement consistency is enforced through posterior-guided sampling without retraining, preserving coherence with the learned data manifold while reducing reconstruction ambiguity and variability.
+These properties make Cas-Sensing particularly suitable for practical scientific and engineering applications.
+In fluid dynamics, reliable reconstruction of coherent flow structures from sparse sensors supports monitoring and control tasks.
+In climate and ocean science, the ability to recover coarse-scale background fields together with refined-scale variability enables robust environmental assessment.
+In geophysical and engineering sensing, generalization across sensor layouts and geometric configurations facilitates deployment in complex, real-world environments.
+Despite these advantages, several limitations remain.
+The current implementation is restricted to two-dimensional spatial domains and does not explicitly incorporate temporal dynamics, limiting applicability to fully spatiotemporal or three-dimensional systems.
+In addition, reliance on a vanilla denoising diffusion probabilistic model introduces practical constraints on sampling efficiency, which may become significant in coarse-scale applications.
+These limitations also suggest promising directions for future research.
+Extending the cascaded sensing paradigm to higher-dimensional settings would enable reconstruction of volumetric fields and complex 3D geometries.
+Incorporating temporal modeling into the cascade could further generalize the framework to spatiotemporal reconstruction and prediction, supporting real-time monitoring and forecasting tasks.
+Moreover, integrating more advanced generative paradigms, such as flow-matching-based or accelerated diffusion methods, may substantially improve sampling efficiency while preserving reconstruction quality.
+Overall, Cas-Sensing represents a general, extensible, and principled framework for data-driven scientific sensing.
+By explicitly structuring uncertainty, separating coarse- and fine-scale reconstruction responsibilities, and modeling conditional distributions rather than deterministic mappings, this work provides a systematic approach to addressing non-uniqueness and multi-scale complexity in sparse inverse problems.
+The proposed paradigm lays the groundwork for future sensing and inference systems that are both robust to severe data incompleteness and adaptable to complex physical environments.
+Acknowledgements
+The authors wish to express their gratitude for the financial support received from Guangdong Provincial Fund - Special Innovation Project (2024KTSCX038); Research Grants Council of Hong Kong through the Research Impact Fund (R5006-23); the Guangdong Provincial Key Lab of Integrated Communication, Sensing and Computation for Ubiquitous Internet of Things (No.
+2023B1212010007).
+The authors also thank Mr.
+Xiao DI for generously providing the circular cylinder flow dataset used in this study.